OpenStudy (anonymous):

solve the following equation for B M=b+c / 2

OpenStudy (anonymous):

\[M = \frac{ b+c }{ 2 }\]

OpenStudy (tkhunny):

Have you considered multiplying by 2?

OpenStudy (anonymous):

\[M=\frac{b+c}{2}\]\[2M=b+c\]\[2M-c=b\]